# The quantum criticality of the Standard Model and the hierarchy problem

### v1

@@ Abstract
The naturalness principle has long guided efforts to understand what lies beyond the Standard Model of particle physics, with the hierarchy problem being a prominent example. In this work, we revisit the impact of quantum corrections on the fine-tuning of the low-energy effective theory and its phase structure. For this purpose, we employ the Wilsonian functional renormalization group which conveniently captures both, logarithmic and polynomial, scalings in dimensionful parameters along the flow and allows for an interpretation in terms of critical phenomena. Within this framework, we show that crucial and scheme-independent information about the infrared Higgs phases and the associated quantum phase transition can be directly extracted. We build on the connection between the hierarchy problem and the near-criticality of the Standard Model, identifying a particularly enhanced tuning, and perform a comprehensive and model-independent analysis of the sensitivity to heavy new physics. Finally, we illustrate the versatility of this framework by performing a simplified search for new physics coupled to the Higgs sector which can soften the sensitivity to high-energy scales, rediscovering also the large-anomalous-dimension solution to the hierarchy problem.
@@ 1 Introduction
So far, the Standard Model (SM) of particle physics has proven to be remarkably successful in describing nature at the most fundamental level accessible to current experimental efforts. Despite this success, established observational evidence indicates that the SM cannot be the ultimate theory of nature but rather provides a highly accurate description of its low-energy limit. The origin of dark matter, neutrino masses, and the baryon asymmetry of the universe remain unexplained within the current framework, demonstrating the need for physics beyond the SM (BSM).
@@
To address these open problems, various extensions of the SM have been proposed. Since there is no evidence that the theoretical formulation of the SM breaks down below the Planck scale, such extensions have also been largely motivated by unresolved puzzles within the SM itself, such as the reason for the hierarchies in fermion masses and mixings or the absence of CP violation in strong interactions. One particularly intriguing puzzle is the origin of the electroweak (EW) scale, determined by the Higgs mass or vacuum expectation value, and its relation to potential new high-energy scales beyond the SM. The strong sensitivity of the infrared (IR) phenomenology to ultraviolet (UV) embeddings, rooted in the Higgs mass parameter, gives rise to the hierarchy problem. Such information is encoded in quantum corrections to correlation functions, in particular, the Higgs two-point function. Assuming naive canonical power counting, dimensionful quantities receive two types of corrections: logarithmic and power law. The large sensitivity to high-energy scales arises in presence of the latter and, specifically, from the competition between both types of corrections. Having to precisely set parameters at high-energy scales to obtain the correct physics in the infrared stands in tension with the principle of naturalness tHooft:1979rat ; Giudice:2008bi ; Dirac:1938mt ; Craig:2022eqo . While this is not an inconsistency that must be resolved in the ultimate fundamental embedding of the SM (unlike, for example, dark matter), the principle of naturalness provides a guiding criterion to constrain viable theoretical scenarios.
@@
To investigate the sensitivity between UV and IR physics, the power-law corrections to the Higgs mass play a crucial role. These corrections are proportional to the square of a given UV scale, indicating the limit of validity of the effective theory. This sensitivity of the IR physics to UV parameters was originally observed in studies of Grand Unification theories (GUTs), where the Higgs mass had to be finely tuned relative to the GUT scale and became generically denoted as the gauge-hierarchy problem Gildener1976 ; Weinberg1979a . In standard perturbative approaches, these corrections only emerge at thresholds because of the mass-independent nature of the regularization. Consequently, to access these power-law contributions with these methods, one must postulate a specific UV completion. In contrast, Wilsonian renormalization schemes Wilson:1974mb , which use a cutoff to progressively integrate quantum fluctuations, naturally account for power-law corrections. In this context, the UV cutoff can be interpreted as the result of integrating out a more fundamental embedding. Importantly, this bottom-up approach does not require specifying a UV completion to manifest this polynomial sensitivity to UV scales.
@@
In this work, we investigate the role of such corrections in the context of quantum criticality and naturalness in the SM using the functional Renormalization Group (fRG) Wetterich:1992yh . The fRG is a quantitative and robust method that has been successfully applied across a wide range of fields, from condensed matter, confining gauge theories, and dynamical chiral symmetry breaking to quantum gravity; see Dupuis:2020fhh for a review. In high-energy particle physics, including the SM and its extensions, the fRG has been used to study individual sectors such as the Higgs Gies:2013fua ; Gies:2014xha ; Eichhorn:2015kea ; Reichert:2017puo ; Sondenheimer:2017jin ; Gies:2017ajd ; Held:2018cxd ; Eichhorn:2020upj , Yukawa Borchardt:2016xju ; Gies:2017zwf ; Ellwanger:1992us , and gauge-Yukawa Gies:2019nij ; Gies:2013pma ; Gies:2018vwk ; Gies:2016kkk ; Gies:2015lia ; Gies:2023jzd sectors. The present work builds on the full implementation of the SM developed in Pastor-Gutierrez:2022nki ; Goertz:2023pvn , which consistently includes all SM sectors from Planckian to sub-QCD scales.
@@
Due to its structural advantage in accessing power-law sensitivity in a convenient manner, the Wilsonian approach has already been used in previous studies Krajewski:2014vea ; Krajewski:2015eea ; Gies:2023jzd . The first study in this direction was performed in Krajewski:2014vea ; Krajewski:2015eea for a Higgs-Yukawa toy model, exploring the fine-tuning required as a function of the new-physics scale and its implications for vacuum stability. More recently, in Gies:2023jzd , the interplay between different sources of chiral symmetry breaking in the SM - through Higgs and emergent condensates - was studied in detail, with a special focus on its impact on fine-tuning. There, it was shown that within the SM class of theories, there exists a natural limit where fine-tuning is minimized as the QCD and EWSB scales become similar. These studies highlight the strength of Wilsonian methods in assessing more natural configurations in theory space.
@@
In this work, we add to these previous studies by analyzing the near-critical nature of the SM, understood as its proximity to an EW-symmetric solution, and its close relation to the hierarchy problem. To this end, we exploit the analogy between the IR cutoff and a physical temperature to investigate the phase structure of the SM class of theories along the RG flow. Furthermore, we study the fine-tuning problem in the SM effective theory from a bottom-up perspective and explore how deformations in theory space can affect it. This allows for a simplified yet systematic tool for the search of new physics based on the principle of naturalness. This is particularly relevant nowadays due to the absence of new physics at scales expected from naturalness principles, a problem commonly known as the little hierarchy problem. This issue has motivated the development of Little Higgs models ArkaniHamed:2001nc ; ArkaniHamed:2002qx ; ArkaniHamed:2002qy ; Low:2002ws ; Ahmed:2023qsm , where the scale of new physics is lifted by an order of magnitude. Similar mechanisms appear in other frameworks, for example with TeV-scale hidden sectors; see, e.g. Chacko:2005pe ; Batell:2022pzc ; Blasi:2019jqc ; Blasi:2020ktl ; Blasi:2022hgi ; Angelescu:2023usv ; Bally:2022naz ; Chung:2023gcm . Ultimately, the challenge of accounting for the observed low-energy phenomena with minimal fine-tuning in the presence of a large hierarchy of scales remains an open question.
@@
This work is organized as follows. First, in Section 2, we study the SM results obtained from functional methods and place them in the context of the standard perturbative picture. This comparison allows us to identify what information is already encoded in the Wilsonian scheme, such as the evolution across different Higgs phases. We then proceed in Section 3 to analyze the phase structure of the SM class of theories and its evolution along the RG flow, focusing on the scalar sector. Having set the stage, in Section 4, we study the quantum phase transition present in the class of theories and its close relation to the hierarchy problem, highlighting the impact of the near-criticality of the SM even when new physics enters at the TeV scale. In Section 5, we quantify the fine-tuning associated with deforming the SM in various directions in parameter space, and use this to search for new physics scenarios that minimize fine-tuning in Section 6. Finally, we conclude in Section 7.
@@ 2 The Standard Model from a functional point of view
In Figure 1 we can identify three different regimes determined by the shape of the effective average potential, namely the effective potential at finite $k$. While the full effective potential is recovered as $k\to 0$, the effective average Higgs potential displays the effective shape at the respective scale and encodes the quantum fluctuations from $\Lambda_{\textrm{UV}}$ to $k$. We extract its shape from the integrated values of $\bar{\mu}^{2}$ and $\bar{\lambda}$ at a given scale. Starting at the Planck scale and flowing down, the curvature mass is positive and the Higgs self-coupling negative, see regime (C). This renders an effective shape of the Higgs potential with a metastable minimum where the Higgs field is found and expected not to tunnel. Moreover, at $k_{\textrm{meta}}\simeq 10^{10}\,$eV, $\bar{\lambda}$ turns positive and the effective shape of the Higgs potential recovers a trivially stable shape (regime (B) in Figure 1). At $k\simeq 5\cdot 10^{3}$ GeV the flow of the curvature mass rapidly changes signs and induces, at $k_{\textrm{SSB,1}}\simeq 950$ GeV, a non-trivial minimum of the potential (see regime (A)). This corresponds to the Higgs’ vacuum expectation value whose dimensionful form freezes at $v_{k\lesssim 200\textrm{GeV}}=246\,$ GeV. This non-trivial vacuum generates the masses of the EW gauge bosons and fermions via the Higgs mechanism Englert:1964et ; Higgs:1964pj ; Guralnik:1964eu . For the cutoff profile of $v$ and all the SM Euclidean masses, see Pastor-Gutierrez:2022nki .
@@ 3 Phase diagram of the Standard Model class of theories
[…]
@@
[…]
@@ 4 Quantum phase transition and criticality
[…]
@@ 5 Quantifying naturalness and the hierarchy problem
Figure 6 highlights two dominant sources of fine-tuning: the magnitude of the new physics scale ($\Lambda_{\textrm{UV}}$) and the distance of the theory from the critical surface. While the former is well known, the latter is often overlooked, as it requires identifying the presence and structure of the critical surface, making it less apparent. However, this second source induces a sensitivity that can exceed the usual quadratic dependence on $\Lambda_{\textrm{UV}}$ in theories near a quantum phase transition, rendering it highly relevant for a proper quantification of the fine-tuning. This dependence is governed by the universal scaling behavior in the vicinity of the massless critical theory. In order to accurately describe the different scalings (particularly with respect to marginal parameters), both power-law and logarithmic corrections have to be taken into account. This is naturally achieved within the present RG approach.
@@
The BG measure, as defined in (19), identifies the most sensitive direction in theory space as the one determining the degree of fine-tuning. In the results shown in Figure 6, we considered the variations only in $\bar{\mu}^{2}$ as the single UV parameter and used $v$ as the only observable IR observable. We now extend the analysis to explore other directions in UV theory space, as well as additional IR observables. In the left panel of Figure 7, we display the mass spectrum of the EW gauge bosons, Higgs, and top quark as a function of relative variations in the curvature mass at $\Lambda_{\textrm{UV}}=10^{10}\,$eV. For theories in close proximity to the SM, the sensitivity of these IR observables is primarily driven by variations in $v$, thereby justifying the earlier approximation of neglecting the variation of other couplings.
@@
We now proceed to investigate the sensitivity of IR observables to variations along other directions in theory space beyond the curvature mass, such as the gauge and Yukawa couplings. From the perspective of the EFT, these variations can be linked to integrating out different types of new physics in the respective sectors. For example, new physics coupled predominantly to the top quark will manifest itself in the quadratic sensitivity via the second term in square brackets in (2). In Figure 8, we present the dimensionless version of the BG measure, where the amount of fine-tuning has been normalized by $(\Delta\Lambda)^{2}$ to remove the leading canonical scaling. The Figure shows the dependence of this normalized measure on relative variations in different directions in theory space. The trajectories shown were initiated at $\Lambda_{\textrm{UV}}=10^{8}\,$eV, and only slight differences are observed for other choices of the UV cutoff scale.

### v2

@@ Abstract
The naturalness principle has long guided efforts to understand physics beyond the Standard Model, with the hierarchy problem as the central issue. We revisit the role of quantum corrections in the fine-tuning of the low-energy effective description and its phase structure. We implement, for the first time in this context, the full Standard Model within the Wilsonian functional renormalization group. Crucially, this method captures conveniently both logarithmic and quadratic scalings, which must both be considered in the tuning, and allows us to provide a new generic and quantitative study of fine-tuning and its interpretation in terms of critical phenomena. We emphasize on the connection between the hierarchy problem and the near-criticality of the Standard Model and extract scheme-independent information on the infrared Higgs phases and the associated quantum phase transition as well as discuss a related enhanced fine-tuning usually not considered in tuning estimates. Finally, we illustrate the framework’s versatility by exploring new physics coupled to the Higgs sector that can soften high-scale sensitivity, recovering also the large-anomalous-dimension solution to the hierarchy problem.
@@ 1 Introduction
So far, the Standard Model (SM) of particle physics has been remarkably successful in describing nature at the most fundamental level accessible to experiment. Despite this success, established evidence shows that the SM cannot be the ultimate theory but rather a highly accurate low-energy limit. It fails to explain the origin of dark matter, neutrino masses, and the baryon asymmetry of the universe, indicating the need for physics beyond the SM (BSM).
@@
Many extensions of the SM have been proposed to address these shortcomings. Since no breakdown of the SM is observed below the Planck scale, such extensions are also motivated by internal puzzles of the SM itself, such as the fermion mass hierarchies or the absence of CP violation in strong interactions. A particularly intriguing puzzle is the origin of the electroweak (EW) scale, set by the Higgs mass or vacuum expectation value, and its relation to possible new high-energy scales. The strong dependence of infrared (IR) physics on ultraviolet (UV) embeddings, rooted in the single dimensionful parameter of the SM (the Higgs mass), constitutes the hierarchy problem. Such sensitivity is encoded in quantum corrections to correlation functions – here the Higgs two-point function. Canonical power counting implies that dimensionful quantities receive two kinds of corrections: logarithmic and power-law. The tension arises in the presence of the latter and, specifically, from the interplay of both types. Obtaining the correct IR physics then requires fine adjustment of UV parameters, in conflict with the principle of naturalness tHooft:1979rat ; Giudice:2008bi ; Dirac:1938mt ; Craig:2022eqo . While not a logical inconsistency that must be solved (unlike, e.g. dark matter), naturalness provides a guiding criterion for viable theoretical scenarios.
In particular, the absence of supersymmetry at current collider energies has renewed interest in alternative approaches to address the hierarchy problem. Experimentally, the lack of new physics up to multi-TeV scales points to a little hierarchy problem, motivating mechanisms that provide a natural extension and raise the expected new-physics scale. Examples include Little Higgs models ArkaniHamed:2001nc ; ArkaniHamed:2002qx ; ArkaniHamed:2002qy ; Low:2002ws ; Ahmed:2023qsm , where the Higgs emerges as a pseudo–Goldstone boson, lifting the scale by roughly an order of magnitude, as well as TeV-scale hidden sectors Chacko:2005pe ; Batell:2022pzc ; Bally:2022naz ; Chung:2023gcm . In this context, considering the little hierarchy as a genuine result of a single more fundamental model, rather than an independent issue, might lead to comprehensive solutions, see e.g. the recent Blasi:2019jqc ; Blasi:2020ktl ; Blasi:2022hgi ; Angelescu:2023usv ; Chung:2023gcm ; Ahmed:2023qsm ; Cacciapaglia:2020kgq ; Cacciapaglia2022a .
@@
For a comprehensive understanding of the problem, which may potentially motivate the study of unexplored scenarios, the power-law corrections to the Higgs mass, being central to the UV–IR sensitivity, need to be analyzed in detail. They scale with the square of a given UV cutoff, which marks the limit of validity of the effective theory. This issue was first highlighted in studies of Grand Unification Theories (GUTs), where the Higgs mass had to be finely tuned relative to the GUT scale, giving rise to the gauge hierarchy problemGildener1976 ; Weinberg1979a . In standard perturbative approaches, these corrections only emerge at thresholds because of the mass-independent nature of the regularization. Consequently, to access these power-law contributions with these methods, one must postulate a specific UV completion. In contrast, Wilsonian renormalization schemes Wilson:1974mb , which use a cutoff to progressively integrate quantum fluctuations, naturally account for power-law corrections. In this context, the UV cutoff can be interpreted as the result of integrating out a more fundamental embedding. This bottom-up approach does not require specifying a UV completion to manifest this polynomial sensitivity to high scales.
@@
In this work, we investigate the role of corrections to the Higgs mass in the context of quantum criticality and naturalness in the SM using the functional Renormalization Group (fRG) Wetterich:1992yh . The fRG is a quantitative and robust method that has been successfully applied across a wide range of fields, from condensed matter, confining gauge theories, and dynamical chiral symmetry breaking to quantum gravity; see Dupuis:2020fhh for a review. In high-energy particle physics, including the SM and its extensions, the default approach relies on perturbative methods. However, the fRG has been used to study individual sectors such as the Higgs Gies:2013fua ; Gies:2014xha ; Eichhorn:2015kea ; Reichert:2017puo ; Sondenheimer:2017jin ; Gies:2017ajd ; Held:2018cxd ; Eichhorn:2020upj , Yukawa Borchardt:2016xju ; Gies:2017zwf ; Ellwanger:1992us , and gauge-Yukawa Gies:2019nij ; Gies:2013pma ; Gies:2018vwk ; Gies:2016kkk ; Gies:2015lia ; Gies:2023jzd sectors. The present work builds on the full implementation of the SM developed in Pastor-Gutierrez:2022nki ; Goertz:2023pvn , which consistently includes all SM sectors from Planckian to sub-QCD scales.
@@
Despite its structural advantage in exposing power-law sensitivity, the Wilsonian approach has only been applied in a few studies of naturalness Krajewski:2014vea ; Krajewski:2015eea ; Gies:2023jzd . The first study in this direction was performed in Krajewski:2014vea ; Krajewski:2015eea for a Higgs-Yukawa toy model, exploring the fine-tuning required as a function of the new-physics scale and its implications for vacuum stability. More recently, in Gies:2023jzd , the interplay between different sources of chiral symmetry breaking in the SM - through Higgs and emergent condensates - was studied in detail, with a special focus on its impact on the quantum phase transition. There, it was shown that within the SM class of theories, there exists a natural limit where fine-tuning is minimized as the QCD and EWSB scales become similar. These studies highlight the strength of Wilsonian methods in assessing more natural configurations in theory space.
@@
Here, we add to previous studies by analyzing the near-critical nature of the SM, understood as its proximity to an EW-symmetric solution, and its relation to the hierarchy problem in the context of new physics. To this end, we exploit the analogy between the IR cutoff and a physical temperature to investigate the complete phase structure of SM-like theories along the RG flow. In addition, we study the fine-tuning problem in the SM effective theory from a bottom-up perspective and explore how BSM deformations in theory space affect it. This provides a simplified but systematic tool for the search for new physics based on the principle of naturalness, which is particularly relevant in light of the absence of new physics at scales expected from naturalness arguments. The following analysis is, to the best of our knowledge, the first fRG-based naturalness study that incorporates the full Standard Model and provides quantitative results.
@@
This work is organized as follows. First, in Section 2, we introduce the SM results obtained from functional methods and place them in the context of the standard perturbative picture. This comparison allows us to identify what information is already encoded in the Wilsonian scheme, such as the evolution across different Higgs phases. We then proceed in Section 3 to analyze the phase structure of the general SM class of theories and its evolution along the RG flow, particularly focusing on the scalar sector. Having set the stage, in Section 4, we study the quantum phase transition present in the class of theories and its close relation to the hierarchy problem, highlighting the impact of the near-criticality of the SM even when new physics enters at the TeV scale, which is particularly important in the context of the little hierarchy problem. In Section 5, we quantify the fine-tuning associated with deforming the SM in various directions in parameter space, and use this to search for new physics scenarios that minimize fine-tuning in Section 6. Finally, we conclude in Section 7.
@@ 2 The Standard Model from a functional point of view
In Figure 1 we can identify three different regimes determined by the shape of the effective average potential, namely the effective potential at finite $k$. While the full effective potential is recovered as $k\to 0$, the effective average Higgs potential displays the effective shape at the respective scale and encodes the quantum fluctuations from $\Lambda_{\textrm{UV}}$ to $k$. We extract its shape from the integrated values of $\bar{\mu}^{2}$ and $\bar{\lambda}$ at a given scale. Starting at the Planck scale and flowing down, the curvature mass is positive and the Higgs self-coupling negative, see regime (C). This renders an effective shape of the Higgs potential with a metastable minimum where the Higgs field is found and expected not to tunnel. Moreover, at $k_{\textrm{meta}}\simeq 10^{10}\,$GeV, $\bar{\lambda}$ turns positive and the effective shape of the Higgs potential recovers a trivially stable shape (regime (B) in Figure 1). At $k\simeq 5\cdot 10^{3}$ GeV the flow of the curvature mass rapidly changes signs and induces, at $k_{\textrm{SSB,1}}\simeq 950$ GeV, a non-trivial minimum of the potential (see regime (A)). This corresponds to the Higgs’ vacuum expectation value whose dimensionful form freezes at $v_{k\lesssim 200\textrm{GeV}}=246\,$ GeV. This non-trivial vacuum generates the masses of the EW gauge bosons and fermions via the Higgs mechanism Englert:1964et ; Higgs:1964pj ; Guralnik:1964eu . For the cutoff profile of $v$ and all the SM Euclidean masses, see Pastor-Gutierrez:2022nki .
@@ 3 Phase diagram of the Standard Model class of theories
$\partial_{t}\bar{\lambda}\left(\bar{\mu}^{*\,2},\bar{\lambda}^{*},\vec{g}\right)=0$ and $\partial_{t}\bar{\mu}^{2}\left(\bar{\mu}^{*\,2},\bar{\lambda}^{*},\vec{g}\right)=0\,$, (12)
@@
The partial fixed point condition in (15) is also commonly known as the massless trajectory given that on this the scalar field never decouples, neither with a positive nor a negative curvature mass. Note that while in the pure scalar case the mass parameter $\bar{\mu}^{2}_{k\to 0}\to\bar{\mu}^{*\,2}=0$, if (3.4) is fulfilled, in SM-like systems $\bar{\mu}^{2}_{k\to 0}\to\bar{\mu}^{*\,2}\neq 0$. Nevertheless, in both cases a non-decoupling of the Higgs scalar is realized, and no explicit scale is generated. We may refer to this particular solution as the trajectory on the critical surface Aoki:2012xs ; Wetterich:2019qzx and it minimizes the impact of the relevant scaling. Although partial fixed points such as the one in (15), do not have to be related to symmetries of the classical action, they usually are. In the present case with the curvature mass, it is clear how the massless condition renders scale symmetry Wetterich:1983bi ; Wetterich:1990an ; Wetterich:2019qzx .
@@ 4 Quantum phase transition and criticality
This illustrates that drastically different IR dynamics—either with or without SSB in the Higgs sector—can emerge from small, natural variations of UV parameters. Theories in the symmetric phase, lacking a Higgs condensate, would exhibit massless fermions and electroweak gauge bosons until chiral symmetry breaking is dynamically triggered by the strong coupling. Consequently, the two non-Abelian gauge groups—SU(3)c and chiral SU(2)L—would confine sequentially, with color confinement occurring first due to the faster running of the strong coupling. In the absence of explicit fermion masses, the strong sector would break the global chiral symmetry $(N_{f})_{R}\times SU$$(N_{f})_{L}$, giving rise to exactly massless Nambu–Goldstone bosons (pions). This condensate would then indirectly generate fermion and EW boson masses, with a spectrum largely independent of flavor. Altogether, the absence of a Higgs $v$ would yield a Universe radically different from the one we observe.
@@ 5 Quantifying naturalness and the hierarchy problem
Figure 6 highlights two dominant sources of fine-tuning: the magnitude of the new physics scale ($\Lambda_{\textrm{UV}}$) and the distance of the theory from the critical surface. While the former is well known, the latter is often overlooked, as it requires identifying the presence and structure of the critical surface, making it less apparent. However, this second source induces a sensitivity that can exceed the usual quadratic dependence on $\Lambda_{\textrm{UV}}$ in theories near a quantum phase transition, rendering it highly relevant for a proper quantification of the fine-tuning, in particular in the context of the little hierarchy problem. This dependence is governed by the universal scaling behavior in the vicinity of the massless critical theory. In order to accurately describe the different scalings (particularly with respect to marginal parameters), both power-law and logarithmic corrections have to be taken into account. This is naturally achieved within the present RG approach.
@@
The BG measure, as defined in (19), identifies the most sensitive direction in theory space as the one determining the degree of fine-tuning. In the results shown in Figure 6, we considered the variations only in $\bar{\mu}^{2}$ as the single UV parameter and used $v$ as the only observable IR observable. We now extend the analysis to explore other directions in UV theory space, as well as additional IR observables. In the left panel of Figure 7, we display the mass spectrum of the EW gauge bosons, Higgs, and top quark as a function of relative variations in the curvature mass at $\Lambda_{\textrm{UV}}=10^{10}\,$GeV. For theories in close proximity to the SM, the sensitivity of these IR observables is primarily driven by variations in $v$, thereby justifying the earlier approximation of neglecting the variation of other couplings.
@@
We now proceed to investigate the sensitivity of IR observables to variations along other directions in theory space beyond the curvature mass, such as the gauge and Yukawa couplings. From the perspective of the EFT, these variations can be linked to integrating out different types of new physics in the respective sectors. For example, new physics coupled predominantly to the top quark will manifest itself in the quadratic sensitivity via the second term in square brackets in (2). In Figure 8, we present the dimensionless version of the BG measure, where the amount of fine-tuning has been normalized by $(\Delta\Lambda)^{2}$ to remove the leading canonical scaling. The Figure shows the dependence of this normalized measure on relative variations in different directions in theory space. The trajectories shown were initiated at $\Lambda_{\textrm{UV}}=10^{8}\,$GeV, and only slight differences are observed for other choices of the UV cutoff scale.